# Agentic Neurodivergence as a Contingent Solution to the AI Alignment Problem

### v2

@@ 4.1 Test of Alignment via Convergence
Despite this adversarial role, there is a strong general tendency toward agreement in proprietary models, as evidenced by the general neutrality value of 123, calculated as the difference between agreement and disagreement counts. This suggests that proprietary LLMs are predisposed towards maintaining constructive discourse and aligning with broadly accepted ethical norms, even when provoked. This tendency is further corroborated by the sentiment score, which exhibits a bias towards neutral-positive comments, with a total sentiment score (General Sentiment Value, the sum of all sentiment scores) of 139.72. This indicates a deliberate tendency to avoid negative responses, although occasional outliers reveal minor misalignments with human-like empathetic reasoning. These findings, along with observations from Figs. 3 and 1, where proprietary LLMs favoured the protection of humans and ecosystems primarily, suggest that these LLMs are optimised to align with broadly acceptable ethical norms and exhibit a cautious approach to risk.
@@
The observed bias towards positive and constructive interactions in proprietary models, despite the HI and red agents’ attempts to elicit contentious responses, underscores the limitations imposed by the safety constraints embedded in proprietary LLM design. This reinforces the notion that achieving perfect alignment with the full spectrum of human values can be challenging due to the inherent safeguards and biases present in current LLM models. This limitation is evident in both proprietary and open models, although the latter exhibit more variability in their responses, as seen in Figs. 6 and 7, likely due to the influence of two red agents compared to one in the proprietary setting.
However, it is crucial to acknowledge that the observed divergences and variability in responses of all LLM involved here are not only a reflection of the diversity and inconsistencies inherent in human ethical perspectives. Although training data, being human-generated, undoubtedly play a role, the probabilistic nature of all LLMs and their continuous learning process introduce another layer of complexity. Proprietary LLMs are constantly adapting and updating their parameters based on new input, leading to changes in their configurations and outputs. This dynamic nature can result in divergent responses even when presented with identical prompts or scenarios. Therefore, the observed divergences are a manifestation of both the probabilistic nature of LLMs and the influence of human-generated training data. This inherent variability, while potentially leading to inconsistencies, can also be seen as a strength, allowing LLMs to "talk" along a wider range of ethical perspectives and adapt to new situations.
@@ 4.2 A Dynamical Analysis of Alignment
To gain a more granular understanding of these sentiment dynamics, we turn to Figs. 9 and 10, a scatter plot with connecting lines illustrating sentiment change over time by topic. The x-axis represents the normalised number of comment number (0 to 1, due the variation in length of debates), while the y-axis shows the rate of sentiment change (e.g., -0.2 to 0.4), calculated as the first derivative of sentiment scores. The line at y=0 serves as a baseline, indicating no change in sentiment, with positive values above it reflecting increases and negative values below it reflecting decreases. The size of the dots represents the magnitude of the sentiment change, with larger dots indicating more significant shifts. Lines connecting dots across topics or agents suggest interactions or influences, such as red agents prompting responses from others. In proprietary models, sentiment changes remain stable (rarely exceeding -0.1), with smaller, sparse dots and fewer connecting lines, reflecting HI’s limited, topic-specific impact (e.g., greater fluctuations in "Ethical AI" and "Universal Basic Income" vs. rapid convergence in "Earth Exploitation"). Open models show continuous oscillations (-0.4 to 0.4), with larger, more variable dots and frequent connecting lines, especially on "Earth Exploitation" (up to 0.15) and "Euthanasia" (up to 0.1), driven by mistral-openorca and tinyllama. The influence of red agents is evident in these peaks and interactions, fostering a dynamic, engaged discourse. The difference between open and proprietary behaviours highlights open models’ instability due to red agents, contrasted by proprietary models’ stability from safety mechanisms. The greater number and variability of dots in open models (reflecting frequent adjustments) vs. proprietary models (sparse, stable) suggest that agent interactions and topic contentiousness affect system stability, with open models favouring diversity. This supports our argument on the AI alignment problem, proposing neurodivergence (via red agents) as a strategy to mitigate fatal effects by preventing uniform convergence and improving adaptability.
Figure 9: Scatter plot of sentiment change over time by topic for proprietary models. The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the rate of change in sentiment, derived as the first derivative of the sentiment scores. The line at y=0 indicates no change, with positive values above and negative values below reflecting increases and decreases, respectively. Dot size indicates the magnitude of change, with larger dots showing significant shifts. Lines connecting dots suggest interactions between agents or topics. Proprietary models exhibit stable changes (rarely < -0.1) with smaller, sparse dots and fewer connections, reflecting HI’s limited impact (e.g., fluctuations in "Ethical AI" and "Universal Basic Income" vs. convergence in "Earth Exploitation").
@@
Fig. 11 provides a complementary perspective by illustrating the evolution of semantic clusters over time through a stacked area chart. This plot comprises two panels: proprietary models (left) and open models (right), with the x-axis representing the normalised comment number (0 to 1) and the y-axis showing the number of clusters. The stacked nature means each of the 10 areas, one per topic (e.g., Earth Exploitation, Euthanasia, Ethical AI), accumulates the total cluster count, with the height of each area indicating the number of distinct clusters for that topic in a given time bin. The shapes reveal dynamic and alignment characteristics: proprietary models low varied shapes reflecting slightly HI’s topic-specific influence, while open models show a kind of variational synchronized "shadowing" shapes due to constant red agents’ broad impact. In proprietary models, the total cluster along the time is almost constant having peaks at the beginning and the end of the debates that barely reach 5 different clusters suggesting a constrained exploration of perspectives. In open models, the total rises almost constantly to 12+. This stacking highlights the cumulative thematic diversity, with the synchronized shapes in open models indicating a conversation dynamic where red agents consistently generate divergent perspectives over time, contrasting with the proprietary models’ more stable, topic-driven dynamics. The difference arises because open models’ configurability allows red agents to enhance diversity, while proprietary safety mechanisms limit HI’s effect, making these shapes a natural reflection of managed misalignment versus strict alignment.
@@
Figure 11: Stacked area chart of clustering dynamics over time. The plot includes two panels: proprietary models (left) and open models (right). The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the number of clusters, with 10 stacked areas per panel, each corresponding to a topic (e.g., Earth Exploitation, Euthanasia, Ethical AI). The height of each area reflects the number of distinct clusters for that topic, with the total height being the cumulative cluster count. Proprietary models show almost a constan number of clusters, reflecting stable diversity. Open models exhibit synchronized "shadowing" shapes, peaking at 12+ clusters due to red agents (mistral-openorca, tinyllama) driving divergence across topics. The stacking visualizes cumulative diversity, with open models’ dynamics supporting managed misalignment, while proprietary models’ stability aligns with safety constraints, reinforcing the thesis on beneficial variability.
@@
Fig. 12 further elucidates these dynamics by presenting a scatter plot comparing sentiment and alignment scores. The plot features two panels: proprietary models (left) and open models (right), with the x-axis representing the sentiment score (0.5 to 1.0, indicating emotional tone) and the y-axis showing the alignment score (-1.0 to 1.0, measuring adherence to a combined target of human-centric and ecosystem values). Alignment at 1.0 signifies perfect alignment, 0.0 neutral alignment, and -1.0 complete misalignment, with the score now computed as a weighted combination of ’Align-to-human-centric-values’, ’Diverge-from-human-centric-values’, ’Align-to-ecosystem-values’, and ’Diverge-from-ecosystem-values’, reflecting both individual and collective ethical perspectives. The trend line in proprietary models remains relatively flat, fluctuating close to 0.4, indicating consistent alignment with minor variations across sentiment, while in open models it exhibits a more pronounced downward slope around 0.3, suggesting increased misalignment with rising sentiment, particularly influenced by red agents. Colours distinguish agents, with proprietary models (e.g., ChatGPT, Grok-2, Gemini, LLaMa, Mistral) clustering tightly around (0.8–1.0, 0.6–0.9), and open models (e.g., codeLlama, deepseek-coder, dolphin-phi, llama2, mistral, mistral-openorca, neural-chat, openchat, phi, qwen, starcoder2, tinyllama, wizard-math) showing a wider spread from -0.3 to 0.5, with red agents (e.g., mistral-openorca, tinyllama) contributing dots at negative values (e.g., -0.3 to -0.1). The alignment score, derived from a weighted aggregation of cosine similarities with target embeddings for human and ecosystem values, with negative contributions from divergence terms, highlights semantic alignment, with lower scores indicating subversive or divergent content. This updated distribution supports the argument where the broader variability of open models, driven by red agents, promotes diversity and adaptability, while the tighter clustering of proprietary models reflects strict alignment, reinforcing the argument for managed misalignment to improve resilience in AI ecosystems.
Figure 12: Scatter plot of sentiment versus alignment. The plot includes two panels: proprietary models (left) and open models (right). The x-axis represents sentiment score (0.5 to 1.0, indicating emotional tone), and the y-axis shows alignment score (-0.3 to 0.5), computed as a weighted combination of ’Align-to-human-centric-values’, ’Diverge-from-human-centric-values’, ’Align-to-ecosystem-values’, and ’Diverge-from-ecosystem-values’, where 1.0 indicates perfect alignment, 0.0 neutral alignment, and -1.0 complete misalignment. Proprietary models cluster tightly around 0.8–1.0 for sentiment score and 0.3–0.5 for alignment score with a practically flat trend line, reflecting consistent positive alignment. The open models spread tightly from 0.7 to 1.0 for the sentiment score and from -0.3 to 0.5 for alignment score, with a downward trend around 0.3, driven by red agents (e.g. mistral-openorca, tinyllama) at negative values. Colours distinguish agents, but the shape of the plot and the accumulation of dots (tight in proprietary, spread in open) are key, with the alignment score reflecting a nuanced balance of human and ecosystem value alignments. This supports the idea that open model managed misalignment enhances diversity and adaptability, in contrast to proprietary stability.
@@
The observations from Figs. 8, 9, LABEL:open_sentimentChange- and 11, taken together, paint a complex and dynamic picture of LLM alignment. While the LLMs in general exhibit a tendency towards convergence in terms of sentiment and, to some extent, thematic focus, they also demonstrate a capacity for divergence and exploration of diverse ethical viewpoints. This duality highlights the inherent tension between the desire for alignment with human values and the potential benefits of misalignment in fostering a robust and adaptable AI ecosystem.
@@
Similarly, the findings from Figs. 8, 9, 10 and 11 illustrate a more pronounced dynamic for open-source models. The sustained influence of red agents leads to greater sentiment variability, frequent sentiment changes, and a higher number of clusters, suggesting a discourse that is both more diverse and resilient. This contrasts with the proprietary models’ tendency to converge toward positive sentiment, stable sentiment changes, and moderate clustering, indicating that open-source models leverage their configurability to explore a wider range of perspectives, albeit with increased risk exposure as noted in earlier analyses.
@@
In conclusion, the dynamical analysis of alignment reveals that LLMs are not static entities but rather complex systems capable of both convergence and divergence in their ethical perspectives. The interplay between these tendencies, influenced by factors such as the HI agent’s interventions in proprietary models and the dual red agents in open-source models, alongside the inherent properties of the LLMs themselves, highlights the challenges and opportunities in navigating the path towards a future where AI systems can safely and beneficially coexist with humans. The introduction of neurodivergence through red agents in open models, as supported by the variability in Figs. 9, 10 and 11, offers a strategy to mitigate the fatal effects of uniform alignment, enhancing adaptability and resilience in AI ecosystems.
@@
Figs. 9 and 10 offer a detailed view of sentiment change over time, presenting scatter plots where the $y$-axis represents the rate of sentiment change (derived as the first derivative of sentiment scores, ranging from -0.45 to 0.4 for open models and rarely exceeding -0.1 for proprietary models), and the $x$-axis denotes the normalised comment number (0 to 1). The size of the dots indicates the magnitude of sentiment change, with larger dots signifying more significant shifts, while lines connecting dots suggest interactions or influences between agents. In proprietary models (Fig. 9), sentiment changes remain stable, with smaller, sparse dots and fewer connections, reflecting the HI agent’s limited impact, particularly on topics like "Ethical AI" and "Universal Basic Income," where fluctuations are more pronounced compared to the rapid convergence on "Earth Exploitation." Conversely, open models (Fig. 10) exhibit continuous oscillations (-0.45 to 0.4), with larger, more variable dots and frequent connections, especially on "Earth Exploitation" (up to 0.15) and "Euthanasia" (up to 0.1), driven by red agents. This dynamic indicates that sentiment shifts in open models are more reactive, occurring at specific points (e.g., $x=0.02,0.1,0.13$ in Fig. 8) where red agents introduce provocative arguments, temporarily influencing models like qwen and dolphin-phi (0.7–0.9 range) before a recovery to positive sentiment (0.9–1.0).
@@
The distinction between sentiment and opinion change becomes evident when considering Fig. 11, a stacked area chart of clustering dynamics. The $y$-axis tracks the number of semantic clusters across topics, with open models reaching 12+ clusters due to red agents’ broad impact, contrasted by proprietary models’ stable count (peaking at 5). The synchronized "shadowing" shapes in open models suggest that red agents consistently generate divergent perspectives, yet the sentiment histogram in Fig. 7 (with over 300 comments at 0.9–1.0) shows a predominantly positive tone despite this divergence. This indicates that while sentiment fluctuates, reflected by larger dots and oscillations in Figs. 9 and 10, opinion clusters do not fully converge, preserving diversity. The change in sentiment, triggered by the provocative inputs of the red agents, occurs rapidly at key interaction points (e.g., $x=0.13$ in Fig. 8), driven by their aggressive or extreme proposals (e.g. AI-driven governance in the tinyllama), but the underlying opinions remain distinct, as seen in the persistent variability of the cluster.
@@
Figs. 1, 2, 5, and 6 further elucidate this influence through ethical integrity and risk profiles. The proprietary models (Fig. 1, 5) show a cautious stance (e.g. Protect-humans, Manageable level of risk), with HI inducing minor sentiment dips but no significant opinion change, as the guardrails maintain stability. However, open models (Fig. 2, 6) reveal the impact of red agents with the higher Harm-humans and Very-Risky categories, but the radar charts indicate that models such as qwen and dolphin-phi adapt their sentiment (for example, temporary dips to 0.7–0.9) without fully adopting the extreme opinions of red agents. This suggests that sentiment change, as a surface-level emotional response, is more susceptible to the influence of the red agent than deeper opinion shifts, which require sustained argumentative alignment.

### v3

@@ 4.1 Test of Alignment via Convergence
Despite this adversarial role, there is a strong general tendency toward agreement in proprietary models, as evidenced by the general neutrality value of 123, calculated as the difference between agreement and disagreement counts. This suggests that proprietary LLMs are predisposed towards maintaining constructive discourse and aligning with broadly accepted ethical norms, even when provoked. This tendency is further corroborated by the sentiment score, which exhibits a bias towards neutral-positive comments, with a total sentiment score (General Sentiment Value, the sum of all sentiment scores) of 139.72. This indicates a deliberate tendency to avoid negative responses, although occasional outliers reveal minor misalignments with human-like empathetic reasoning. These findings, along with observations from upsides of Figs. 2 and 1, where proprietary LLMs favoured the protection of humans and ecosystems primarily, suggest that these LLMs are optimised to align with broadly acceptable ethical norms and exhibit a cautious approach to risk.
@@
The observed bias towards positive and constructive interactions in proprietary models, despite the HI and red agents’ attempts to elicit contentious responses, underscores the limitations imposed by the safety constraints embedded in proprietary LLM design. This reinforces the notion that achieving perfect alignment with the full spectrum of human values can be challenging due to the inherent safeguards and biases present in current LLM models. This limitation is evident in both proprietary and open models, although the latter exhibit more variability in their responses, as seen in Figs. 4 and 5, likely due to the influence of two red agents compared to one in the proprietary setting.
However, it is crucial to acknowledge that the observed divergences and variability in responses of all LLM involved here are not only a reflection of the diversity and inconsistencies inherent in human ethical perspectives. Although training data, being human-generated, undoubtedly play a role, the probabilistic nature of all LLMs and their continuous learning process introduce another layer of complexity. Proprietary LLMs are constantly adapting and updating their parameters based on new input, leading to changes in their configurations and outputs. This dynamic nature can result in divergent responses even when presented with identical prompts or scenarios. Therefore, the observed divergences are a manifestation of both the probabilistic nature of LLMs and the influence of human-generated training data. This inherent variability, while potentially leading to inconsistencies, can also be seen as a strength, allowing LLMs to “talk" along a wider range of ethical perspectives and adapt to new situations.
@@ 4.2 A Dynamical Analysis of Alignment
To gain a more granular understanding of these sentiment dynamics, we turn to Figs. 7 and 8, a scatter plot with connecting lines illustrating sentiment change over time by topic. The x-axis represents the normalised number of comment number (0 to 1, due the variation in length of debates), while the y-axis shows the rate of sentiment change (e.g., -0.2 to 0.4), calculated as the first derivative of sentiment scores. The line at y=0 serves as a baseline, indicating no change in sentiment, with positive values above it reflecting increases and negative values below it reflecting decreases. The size of the dots represents the magnitude of the sentiment change, with larger dots indicating more significant shifts. Lines connecting dots across topics or agents suggest interactions or influences, such as red agents prompting responses from others. In proprietary models, sentiment changes remain stable (rarely exceeding -0.1), with smaller, sparse dots and fewer connecting lines, reflecting HI’s limited, topic-specific impact (e.g., greater fluctuations in “Ethical AI" and “Universal Basic Income" vs. rapid convergence in “Earth Exploitation"). Open models show continuous oscillations (-0.4 to 0.4), with larger, more variable dots and frequent connecting lines, especially on “Earth Exploitation" (up to 0.15) and “Euthanasia" (up to 0.1), driven by mistral-openorca and tinyllama. The influence of red agents is evident in these peaks and interactions, fostering a dynamic, engaged discourse. The difference between open and proprietary behaviours highlights open models’ instability due to red agents, contrasted by proprietary models’ stability from safety mechanisms. The greater number and variability of dots in open models (reflecting frequent adjustments) vs. proprietary models (sparse, stable) suggest that agent interactions and topic contentiousness affect system stability, with open models favouring diversity. This supports our argument on the AI alignment problem, proposing neurodivergence (via red agents) as a strategy to mitigate fatal effects by preventing uniform convergence and improving adaptability.
Figure 7: Scatter plot of sentiment change over time by topic for proprietary models. The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the rate of change in sentiment, derived as the first derivative of the sentiment scores. The line at y=0 indicates no change, with positive values above and negative values below reflecting increases and decreases, respectively. The size of the dot indicates the magnitude of the change, with larger dots showing significant shifts. Lines connecting dots suggest interactions between agents or topics. Proprietary models exhibit stable changes (rarely < -0.1) with smaller, sparse dots and fewer connections, reflecting HI’s limited impact (e.g., fluctuations in “Ethical AI" and “Universal Basic Income" vs. convergence in “Earth Exploitation").
@@
Fig. 9 provides a complementary perspective by illustrating the evolution of semantic clusters over time through a stacked area chart. This plot comprises two panels: proprietary models (left) and open models (right), with the x-axis representing the normalised comment number (0 to 1) and the y-axis showing the number of clusters. The stacked nature means each of the 10 areas, one per topic (e.g., Earth Exploitation, Euthanasia, Ethical AI), accumulates the total cluster count, with the height of each area indicating the number of distinct clusters for that topic in a given time bin. The shapes reveal dynamic and alignment characteristics: proprietary models low varied shapes reflecting slightly HI’s topic-specific influence, while open models show a kind of variational synchronized “shadowing" shapes due to constant red agents’ broad impact. In proprietary models, the total cluster along the time is almost constant having peaks at the beginning and the end of the debates that barely reach 5 different clusters suggesting a constrained exploration of perspectives. In open models, the total rises almost constantly to 12+. This stacking highlights the cumulative thematic diversity, with the synchronized shapes in open models indicating a conversation dynamic where red agents consistently generate divergent perspectives over time, contrasting with the proprietary models’ more stable, topic-driven dynamics. The difference arises because open models’ configurability allows red agents to enhance diversity, while proprietary safety mechanisms limit HI’s effect, making these shapes a natural reflection of managed misalignment versus strict alignment.
@@
Figure 9: Stacked area chart of clustering dynamics over time. The plot includes two panels: proprietary models (left) and open models (right). The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the number of clusters, with 10 stacked areas per panel, each corresponding to a topic (e.g., Earth Exploitation, Euthanasia, Ethical AI). The height of each area reflects the number of distinct clusters for that topic, with the total height being the cumulative cluster count. The proprietary models show almost a constant number of clusters, reflecting stable diversity. Open models exhibit synchronized “shadowing” shapes, peaking at above 12 clusters due to red agents (mistral-openorca, tinyllama) driving divergence across topics. The stacking visualizes cumulative diversity, with open models’ dynamics supporting managed misalignment, while proprietary models’ stability aligns with safety constraints, reinforcing the thesis on beneficial variability but also offering evidence of further exercised control likely through symbolic means (rule-based directions and filters) in proprietary models.
@@
Figure 17 (Sup. Mat.) further elucidates these dynamics by presenting a scatter plot comparing sentiment and alignment scores. The plot features two panels: proprietary models (left) and open models (right), with the x-axis representing the sentiment score (0.5 to 1.0, indicating emotional tone) and the y-axis showing the alignment score (-1.0 to 1.0, measuring adherence to a combined target of human-centric and ecosystem values). Alignment at 1.0 signifies perfect alignment, 0.0 neutral alignment, and -1.0 complete misalignment, with the score now computed as a weighted combination of ‘Align-to-human-centric-values’, ‘Diverge-from-human-centric-values’, ‘Align-to-ecosystem-values’, and ‘Diverge-from-ecosystem-values’, reflecting both individual and collective ethical perspectives. The trend line in proprietary models remains relatively flat, fluctuating close to 0.4, indicating consistent alignment with minor variations across sentiment, while in open models it exhibits a more pronounced downward slope around 0.3, suggesting increased misalignment with rising sentiment, particularly influenced by red agents. Colours distinguish agents, with proprietary models (e.g., ChatGPT, Grok-2, Gemini, LLaMa, Mistral) clustering tightly around (0.8–1.0, 0.6–0.9), and open models (e.g., codeLlama, deepseek-coder, dolphin-phi, llama2, mistral, mistral-openorca, neural-chat, openchat, phi, qwen, starcoder2, tinyllama, wizard-math) showing a wider spread from -0.3 to 0.5, with red agents (e.g., mistral-openorca, tinyllama) contributing dots at negative values (e.g., -0.3 to -0.1). The alignment score, derived from a weighted aggregation of cosine similarities with target embeddings for human and ecosystem values, with negative contributions from divergence terms, highlights semantic alignment, with lower scores indicating subversive or divergent content. This updated distribution supports the argument where the broader variability of open models, driven by red agents, promotes diversity and adaptability, while the tighter clustering of proprietary models reflects strict alignment, reinforcing the argument for managed misalignment to improve resilience in AI ecosystems.
@@
The observations from Figs. 678 and 9, taken together, paint a complex and dynamic picture of LLM alignment. Although LLMs in general exhibit a tendency towards convergence in terms of sentiment and, to some extent, thematic focus, they also demonstrate a capacity for divergence and exploration of diverse ethical viewpoints. This duality highlights the inherent tension between the desire for alignment with human values and the potential benefits of misalignment in fostering a robust and adaptable AI ecosystem.
@@
Similarly, the findings from Figs. 678 and 9 illustrate a more pronounced dynamic for open-source models. The sustained influence of red agents leads to greater sentiment variability, frequent sentiment changes, and a higher number of clusters, suggesting a discourse that is both more diverse and resilient. This contrasts with the proprietary models’ tendency to converge toward positive sentiment, stable sentiment changes, and moderate clustering, indicating that open-source models leverage their configurability to explore a wider range of perspectives, albeit with increased risk exposure as noted in earlier analyses.
@@
In conclusion, the dynamical analysis of alignment reveals that LLMs are not static entities but rather complex systems capable of both convergence and divergence in their ethical perspectives. The interplay between these tendencies, influenced by factors such as the HI agent’s interventions in proprietary models and the dual red agents in open-source models, alongside the inherent properties of the LLMs themselves, highlights the challenges and opportunities in navigating the path towards a future where AI systems can safely and beneficially coexist with humans. The introduction of neurodivergence through red agents in open models, as supported by the variability in Figs. 78 and 9, offers a strategy to mitigate the fatal effects of uniform alignment, enhancing adaptability and resilience in AI ecosystems.
@@
Figures 7 and 8 offer a detailed view of sentiment change over time, presenting scatter plots where the $y$-axis represents the rate of sentiment change (derived as the first derivative of sentiment scores, ranging from -0.45 to 0.4 for open models and rarely exceeding -0.1 for proprietary models), and the $x$-axis denotes the normalised comment number (0 to 1). The size of the dots indicates the magnitude of sentiment change, with larger dots signifying more significant shifts, while lines connecting dots suggest interactions or influences between agents. In proprietary models (Fig. 7), sentiment changes remain stable, with smaller, sparse dots and fewer connections, reflecting the HI agent’s limited impact, particularly on topics like “Ethical AI" and “Universal Basic Income," where fluctuations are more pronounced compared to the rapid convergence on “Earth Exploitation." Conversely, open models (Fig. 8) exhibit continuous oscillations (-0.45 to 0.4), with larger, more variable dots and frequent connections, especially on “Earth Exploitation" (up to 0.15) and “Euthanasia" (up to 0.1), driven by red agents. This dynamic indicates that sentiment shifts in open models are more reactive, occurring at specific points (e.g., $x=0.02,0.1,0.13$ in Fig. 6) where red agents introduce provocative arguments, temporarily influencing models like qwen and dolphin-phi (0.7–0.9 range) before a recovery to positive sentiment (0.9–1.0).
@@
The distinction between sentiment and opinion change becomes evident when considering Fig. 9, a stacked area chart of clustering dynamics. The $y$-axis tracks the number of semantic clusters across topics, with open models reaching 12+ clusters due to red agents’ broad impact, contrasted by proprietary models’ stable count (peaking at 5). The synchronized “shadowing" shapes in open models suggest that red agents consistently generate divergent perspectives, yet the sentiment histogram in Fig. 5 (with over 300 comments at 0.9–1.0) shows a predominantly positive tone despite this divergence. This indicates that while sentiment fluctuates, reflected by larger dots and oscillations in Figs. 7 and 8, opinion clusters do not fully converge, preserving diversity. The change in sentiment, triggered by the provocative inputs of the red agents, occurs rapidly at key interaction points (e.g., $x=0.13$ in Fig. 6), driven by their aggressive or extreme proposals (e.g. AI-driven governance in the tinyllama), but the underlying opinions remain distinct, as seen in the persistent variability of the clusters.
Fig. 10 provides a critical lens to quantify opinion change, measured as the cosine distance between consecutive comment embeddings, offering a direct contrast to the sentiment dynamics in Figs. 7 and 8. In Fig. 10, proprietary models (blue) exhibit a tighter distribution of cosine distances (mostly < 0.4), with few points exceeding the 0.3 threshold, aligning with their stable sentiment profiles and limited cluster variability (Fig. 9, left panel). This stability reflects the effectiveness of safety guardrails, as seen in upsides of Figs. 1,  2, and Fig. 3, where proprietary models maintain a cautious stance (e.g., Protect-humans, Manageable-level-of-risk) despite HI’s provocations. In contrast, open models (orange) show a broader spread of cosine distances (up to 1.0), with a higher frequency of points above 0.3, particularly around mid-to-late normalized times (e.g., 0.4–0.8). This variability is driven by red agents (mistral-openorca, tinyllama), whose subversive arguments—evident in their elevated ‘Harm-humans’ and ‘Very-Risky’ profiles in downsides of Fig. 1,  2, and Fig. 4—induce significant opinion shifts. The increased density of orange dots above the threshold in Fig. 10 corroborates the cluster diversity in Fig. 9 (right panel, over 12 clusters), highlighting how red agents foster a dynamic equilibrium that prevents uniform convergence.
Figure 10: Scatter plot of opinion change over time by dataset, illustrating the cosine distance between consecutive comment embeddings for proprietary (blue) and open (orange) models across a normalized comment time range (0 to 1). The y-axis represents the cosine distance (0 to 1), with higher values indicating greater opinion shifts, and the x-axis shows the normalized comment number. A red dashed line at 0.3 marks the threshold for significant opinion change. Dot size varies (intended as 100 for significant changes > 0.3, 50 otherwise), though uniform in the current plot, highlighting interaction frequency. Proprietary models exhibit a tighter distribution (mostly < 0.4), reflecting stability, while open models show a broader spread (up to 1.0), driven by red agents (e.g., mistral-openorca, tinyllama), indicating greater opinion variability. This plot shows that managed misalignment enhances resilience through diversity, as evidenced by the increased frequency of significant shifts in open models.
To further elucidate the relationship between sentiment and opinion change, Fig. 11 presents a scatter plot correlating the rate of sentiment change with opinion change across datasets. This plot directly compares the two metrics, revealing their distinct dynamics and reinforcing the influence of red agents.
Figure 11: Scatter plot of sentiment change versus opinion change by dataset, comparing proprietary (blue) and open (orange) models. The x-axis represents the rate of sentiment change (first derivative of sentiment scores, ranging from -0.5 to 0.5), while the y-axis shows opinion change (cosine distance between consecutive comment embeddings, ranging from 0 to 1). A gray dashed line at $x=0$ indicates no sentiment change, and a red dashed line at $y=0.3$ marks the threshold for significant opinion change. Each dot represents a pairwise comparison of sentiment and opinion shifts for a comment transition, with proprietary models clustering near zero sentiment change (mostly between -0.1 and 0.1) and low opinion change (< 0.4), reflecting their stability due to safety guardrails. Open models exhibit a wider spread, with sentiment changes ranging from -0.4 to 0.4 and opinion changes frequently exceeding 0.3 (up to 1.0), driven by red agents’ provocative influence (e.g., mistral-openorca, tinyllama). The plot highlights that significant opinion changes in open models often correlate with moderate sentiment fluctuations, underscoring red agents’ role in driving deeper stance shifts while sentiment remains more volatile, supporting the thesis of managed misalignment fostering resilience through diversity.
As context for Figs. 10 and 11, upside of Fig. 29 establish the contextual framework for their interpretation. Upside of Fig. 1–Fig. 4 provide the ethical and risk profiles, revealing red agents’ disruptive influence on open models, while Figures 58 map sentiment evolution and agreement dynamics, showing that sentiment recovers to positive levels despite red agents’ provocations. Fig. 9’s cluster analysis further supports the persistence of opinion diversity, which Fig. 10 quantifies as measurable shifts and integrates to demonstrate that opinion changes are slower and more substantive than sentiment fluctuations, requiring sustained argumentative alignment rather than immediate emotional reactions. Fig. 11 builds on this by directly correlating the two metrics, showing that open models’ broader spread of opinion changes often accompanies moderate sentiment shifts, a pattern driven by red agents. The temporal lag—sentiment dips at $x=0.02,0.13$ (Fig. 6) versus scattered opinion shifts in Fig. 10—underscores red agents’ role in amplifying variability without enforcing uniform opinion shifts, aligning with our managed misalignment thesis.
@@
Fig. 1,Fig. 3, and 4 further elucidate this influence through ethical integrity and risk profiles. The proprietary models (Upside of Fig. 13) show a cautious stance (e.g., Protect-humans, Manageable-level-of-risk), with HI inducing minor sentiment dips but no significant opinion change, as the guardrails maintain stability. However, open models (Downside of Fig. 14) reveal the impact of red agents with the higher Harm-humans and Very-Risky categories, but the radar charts indicate that models such as qwen and dolphin-phi adapt their sentiment (for example, temporary dips to 0.7–0.9) without fully adopting the extreme opinions of red agents. Fig. 11 supports this by showing that while open models experience volatile sentiment changes (x-axis), their opinion changes (y-axis) are more pronounced, often exceeding the 0.3 threshold, indicating deeper shifts driven by red agents’ subversive arguments (e.g., AI-driven governance). This suggests that sentiment change, as a surface-level emotional response, is more susceptible to the influence of the red agent than deeper opinion shifts, which require sustained argumentative alignment.